OpenStudy (styxer):

The maximal value of f(x) = sin x - sqrt(3)*cos(x) (0º<=x<360º) is (?)

OpenStudy (styxer):

I used: sqrt(3) as tg 60º , then sin 60º/cos 60º After a few steps, I got : f(x) = sin(x-60) ---------- cos 60

OpenStudy (styxer):

It's: \[\sin x - \sqrt{3} * \cos x\]

OpenStudy (jdoe0001):

\(\bf f(x)=sin(x)-\sqrt{3}cos(x)\) right?

OpenStudy (jdoe0001):

I assume this is calculus, and thus you're meant to take the derivative of it, and find the maxima?

OpenStudy (styxer):

@jdoe0001 I tried to do that, but I got stuck after taking the derivative :(

OpenStudy (jdoe0001):

alrite, hold the mayo for a few seconds

OpenStudy (royalranger):

Difference rule and the product rule.

OpenStudy (jdoe0001):

\(\bf f(x)=sin(x)-\sqrt{3}cos(x) \\ \quad \\ \cfrac{dy}{dx}=cos(x)-[\sqrt{3}\cdot -sin(x)]\implies cos(x)+\sqrt{3}sin(x) \\ \quad \\ \textit{setting the derivative to 0} \\ \quad \\ 0=cos(x)+\sqrt{3}sin(x)\implies -\sqrt{3}sin(x)=cos(x) \\ \quad \\ \cfrac{-\sqrt{3}sin(x)}{cos(x)}=1\implies \cfrac{sin(x)}{cos(x)}=-\cfrac{1}{\sqrt{3}}\implies tan(x)=-\cfrac{1}{\sqrt{3}} \\ \quad \\ x=tan^{-1}\left( -\cfrac{1}{\sqrt{3}} \right)\)

OpenStudy (jdoe0001):

so... the tangent of it is negative, meaning either, the sine(y) or the cosine(x) is negative on those quadrants and the only two quadrants that happens is II and IV :)

OpenStudy (irishboy123):

\[y = \sin x - \sqrt{3} \cos(x)\] \[ = 2 \left( \frac{1}{2}\sin x - \frac{\sqrt{3}}{2} \cos x \right) \] \[ = 2 \left( \sin \frac{\pi}{6} \sin x - \cos \frac{\pi}{6} \cos x \right) \] \[ = - 2 \cos ( x + \frac{\pi}{6}) \] see if that makes sense, could easily be a typo in there but that's an idea outside of calculus . then think about how a sinusioid oscillates :)

OpenStudy (styxer):

Where this come from ?\[\cfrac{-\sqrt{3}\sin(x)}{\cos(x)}=1\]

OpenStudy (styxer):

@IrishBoy123 hmm... So this equation goes from -2 until 2, right? I can see when x = 150º , -2*cos(180º) = -2 * (-1) = 2 Is this your idea? 2 is maximum, when x = 150º

OpenStudy (royalranger):

I got: \[\frac{ d }{ dx } =\sqrt{3}sinx+cosx\]

OpenStudy (irishboy123):

yes. :)

OpenStudy (styxer):

I see. Thank you very much!

OpenStudy (irishboy123):

you can use calculus but you also get smart :-))

OpenStudy (styxer):

I think I can use the same thought in the expression I've found at the beginning: f(x) = sin(x-60) --------- cos 60 If we want the maximum, sin(of something here) must be 1. So 1/(1/2) will result in 2 as well.

OpenStudy (jdoe0001):

hmmm, stilll wondering where \(\cfrac{-\sqrt{3}\sin(x)}{\cos(x)}=1\) comes from?

OpenStudy (irishboy123):

yes. that checks out. would never have seen it that way :)

OpenStudy (styxer):

@jdoe0001 yes... Is this a property?